### v2

# Batayan: A Filipino NLP benchmark for evaluating Large Language Models

Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities on widely benchmarked high-resource languages. However, linguistic nuances of under-resourced languages remain unexplored. We introduce Batayan, a holistic Filipino benchmark that systematically evaluates LLMs across three key natural language processing (NLP) competencies: understanding, reasoning, and generation. Batayan consolidates eight tasks, three of which have not existed prior for Filipino corpora, covering both Tagalog and code-switched Taglish utterances. Our rigorous, native-speaker-driven adaptation and validation processes ensures fluency and authenticity to the complex morphological and syntactic structures of Filipino, alleviating the pervasive translationese bias in existing Filipino corpora. We report empirical results on a variety of open-source and commercial LLMs, highlighting significant performance gaps that signal the under-representation of Filipino in pre-training corpora, the unique hurdles in modeling Filipino’s rich morphology and construction, and the importance of explicit Filipino language support. Moreover, we discuss the practical challenges encountered in dataset construction and propose principled solutions for building culturally and linguistically-faithful resources in under-represented languages. We also provide a public evaluation suite as a clear foundation for iterative, community-driven progress in Filipino NLP.
[…]
Batayan: A Filipino NLP benchmark for evaluating
Large Language Models
Jann Railey Montalan1,2, Jimson Paulo Layacan3, David Demitri Africa4, Richell Isaiah Flores3, Michael T. Lopez II3, Theresa Denise Magsajo, Anjanette Cayabyab, William Chandra Tjhi1,2 1AI Singapore, 2National University of Singapore, 3Ateneo de Manila University, 4University of Cambridge Correspondence: railey@aisingapore.org
1 Introduction
Figure 1: Batayan is a benchmark that holistically evaluates LLM capabilities on a wide range of Filipino language tasks. The rigorous curation and adaptation done by native Filipinos preserves the complexity and authenticity of Filipino language usage today.
Spurred on by recent advances in computing power, big data, and machine learning, LLMs have come into widespread use due to the emergence of a variety of novel and useful capabilities at scale Hadi et al. (2023). These capabilities have made user applications based on LLMs some of the fastest-growing consumer applications in human history, but have also rendered previous benchmarks insufficiently difficult and diverse Wey (2024); Yang et al. (2023).
Benchmarks are standard datasets used to measure and compare the performance of models against one another. In particular, the increasingly general capabilities of LLMs have necessitated holistic benchmarks which test a diversity of metrics like fairness, truthfulness, and robustness, as well as a variety of tasks like text summarization, casual reasoning, and translation Yang et al. (2023); Guo et al. (2023); Liu et al. (2023). The vast majority of LLM benchmarks evaluate tasks in English, with non-English works being much fewer Liu et al. (2021); Son et al. (2024).
Filipino, despite being the national language of the Philippines and being spoken by over 80 million people, remains an under-resourced language. It is under-represented in multilingual LLMs, and existing corpora are domain-specific, non-multilingual Cajote et al. (2024), and created by non-native speakers Quakenbush (2005); Dita et al. (2009). Filipino is often excluded from multilingual LLM benchmarks, suffer from limited task diversity, or have serious deficiencies in grammatical correctness, completeness, and diversity Bandarkar et al. (2024).
Moreover, Filipino is a complex language which exhibits a highly complex linguistic structure, particularly in its rich morphological system Ramos (2021); Go and Nocon (2017). Its agglutinative nature allows for extensive use of affixation and creating nuance through prefixes, infixes, suffixes, and circumfixes Archibald and O’Grady (2001); Jubilado (2004). These affixes, combined with root words, allow intricate verb conjugations for marking tense, focus, and mood Zamar (2022). Filipino incorporates elements from a wide array of linguistic influences, such as Spanish Bowen (1971); Wolff (2001), Chinese Gonzales (2022); Reid (2018); Chan-Yap (1980), and Malay Wardana et al. (2022); Baklanova (2017), with codeswitching between English and Filipino being common Bautista (1991, 2004).
Our threefold contributions attempt to address these challenges. First, we present Batayan, [^1] The word “batayan” means “basis” or “ground truth”. a holistic Filipino benchmark for evaluating LLMs across 8 distinct natural language processing (NLP) tasks spanning understanding, reasoning, and generation. We place a rigorous emphasis on authenticity to natural Filipino language use through native speaker translation and annotation, addressing translationese bias and other limitations of existing Filipino datasets. Batayan is released as part of SEA-HELM, a publicly available evaluation suite [^2] https://github.com/aisingapore/sea-helm and leaderboard [^3] https://leaderboard.sea-lion.ai for assessing LLMs across linguistic and reasoning tasks for Southeast Asian languages. Second, we provide extensive evaluation results from a comprehensive set of open-source and commercial LLMs on Batayan, from differing sizes (7B–671B parameters) and Filipino-language representation, revealing significant disparities in model performance across different linguistic capabilities in Filipino. Third, we document systematic challenges and methodological considerations in creating high-quality Filipino NLP datasets, particularly highlighting issues in translation fluency, vocabulary adaptation, and the preservation of Filipino’s rich morphological features. These insights provide a practical framework for future development of Filipino language resources.
2 Considerations for Filipino Evaluations
Comp. Task Dataset # test samples Target1 Language Source2 Adaptation Our contribution
NLU PI PAWS Zhang et al. (2019) 2,000 label (2) English English-sourced Native translation
QA Belebele Bandarkar et al. (2024) 900 span (4) Filipino English-adapted Native re-translation
SA PH Elections Cabasag et al. (2019) 5,160 label (3) Taglish Natively-sourced No adaptation
TD PH Elections Cabasag et al. (2019) 5,160 label (2) Taglish Natively-sourced No adaptation
NLR CR Balanced COPA Kavumba et al. (2019) 500 label (2) English English-sourced Native translation
NLI XNLI Conneau et al. (2018) 5,010 label (3) English English-sourced Native translation
NLG AS XL-Sum Hasan et al. (2021) 11,535 summary English English-sourced Native translation
MT FLORES 200 NLLB Team et al. (2022) 1,012 translation Filipino English-adapted Native re-translation
  • 1
    Number of options are shown in parenthesis
  • 2
    English-adapted: previously translated from English; English-sourced: originally in English; Natively-sourced: originally in Taglish
Table 1: Source datasets for each task in Batayan.
2.1 Language of Evaluation
While Filipino is the official language of the Philippines, it has been argued in literature that the de facto lingua franca is Taglish, the practice of code-switching between English and Tagalog Go and Gustilo (2013). Naturally-occurring datasets in Filipino (e.g., mined from social media, recorded interviews) typically contain some code-switching Bautista (2004). We defer the problematization of the debated difference between Tagalog and Filipino, and use the two terms interchangeably.
Filipino, while text-rich, is considered under-resourced, lacking the linguistic data, tools, and resources for effective natural language processing Miranda (2023b). To address data scarcity, developers take advantage of high-resource languages such as English through translation Goyal et al. (2022); Doddapaneni et al. (2023). Furthermore, leveraging multilingual datasets is potent not only because it enhances the performance of models trained on limited resources, but also because code-switching and bilingualism is a common phenomenon in the Philippines Tupas and Martin (2017).
2.2 Language Design Principles
The design of Batayan emphasized linguistic authenticity and representativeness in word choice, sentence structure, and grammar. In terms of language use and word choice, we employed common Filipino and Taglish vocabulary, balancing loanwords with native Filipino terms by prioritizing colloquial usage. This was also done when words had both English and Filipino variants, taking into account spelling and orthographic preferences prevalent among native speakers.
Regarding sentence structure, we prioritized sentences with attention to natural syntax and the choice of ayos (sentence arrangement), namely direct (karaniwang ayos, KA, lit. usual order) or inverted (di-karaniwang ayos, DKA, lit. unusual order) forms Tanawan et al. (2008). In KA, Filipino constructions follow the typical predicate-initial word order Malicsi (2013). In contrast, DKA, which is also referred to linguistically as an ay-inversion, is a type of construction in Filipino where non-predicative constituents (such as simuno, lit. grammatical subject) are “fronted” or shifted to precede the predicate, marking them as the “topic” of the sentence Kroeger (1993). Inverted constructions are often used in formal settings, and could therefore be deemed unnatural in most situations Pizarro-Guevara (2010). This characteristic of Filipino is of interest due to our observation of machine translations preferring DKA (see Section 4.1); hence, the further need for native re-translations with a preference for KA.
2.3 Related Datasets and Benchmarks
Previous work on Filipino language model evaluation has been largely fragmented. Researchers have made efforts in developing datasets for Filipino language tasks such as fake news detection Cruz et al. (2020), named entity recognition Miranda (2023a), natural language inference Cruz et al. (2021), sentiment analysis Villavicencio et al. (2021), and others. However, the field lacks a unified, comprehensive benchmark for systematic evaluation of LLMs across different linguistic dimensions in Filipino.
Unlike existing Southeast Asian benchmarks that either overlook Filipino or treat it through machine translation, Batayan is the first holistic, native-speaker-driven evaluation suite tailored specifically for Filipino NLP. For instance, BHASA provides a multilingual SEA benchmark covering Indonesian, Tamil, Thai, and Vietnamese, but it does not include Filipino Leong et al. (2023). Similarly, IndoNLU focuses exclusively on Indonesian understanding tasks and lacks any generative or reasoning evaluations Wilie et al. (2020). SeaExam and SeaBench introduce local exam and open-ended Vietnamese, Thai, and Indonesian queries but exclude Filipino entirely Liu et al. (2025), and SailCompass evaluates robustness for SEA languages without any Filipino components Guo et al. (2024). SeaEval provides a broad multilingual and multicultural benchmark across 29 datasets for Southeast Asian languages, including Filipino, but still relies primarily on machine translations and generic multilingual prompts Wang et al. (2024).
In contrast, Batayan integrates eight tasks—including three novel Filipino-centric ones—grounded in native translations and rigorous acceptability judgments, ensuring fluency, authenticity, and sensitivity to Filipino’s complex morphological and code-switching patterns.
3 Task and Dataset Curation
3.1 Task Selection
Through BATAYAN, we significantly expand the scope of Filipino LLM evaluation by introducing tasks that assess a gamut of NLP competencies: understanding (NLU), generation (NLG), and reasoning (NLR). Our Filipino benchmark incorporates machine translation (MT), natural language inference (NLI), question answering (QA), sentiment analysis (SA), and toxicity detection (TD). We also release three novel tasks that have not been previously explored in the Filipino context, namely abstractive summarization (AS), causal reasoning (CR), and paraphrase identification (PI).
Our benchmark design and selection of tasks are informed by established multilingual evaluation frameworks, particularly BHASA Leong et al. (2023) and XTREME Hu et al. (2020).
3.2 Dataset Selection
To construct Batayan, we curated open-source corpora with clear provenance when possible, prioritizing datasets that exhibit authentic language use across various domains such as social media, news articles, and other publicly available texts. We identified existing Filipino-language datasets for MT, QA, SA, and TD. Meanwhile, the AS, CR, NLI, and PI tasks lacked an expert-annotated, natively-sourced Filipino corpus. Given the scarcity of high-quality Filipino corpora for these tasks, we chose to natively translate and adapt existing English corpora into Filipino following the methods detailed in Section 3.3. In translation, we deliberately “naturalized” the language use by preferring KA, colloquial terms, and other adaptations to preserve native Filipino speech.
Table 1 outlines key attributes of each dataset, such as output type, domain, language, source, and any adaptations that we applied to ensure alignment with natural Filipino language construction. More detailed statistics can be found in Appendix A.
3.3 Dataset Adaptation and Annotation
Figure 2: Dataset curation process for Batayan.
For each task, entries were filtered by length (20–2000 characters), then randomly sampled $5n$ entries for a target size of $n$, maintaining balanced class distributions.
Native speaker acceptability judgments for quality filtering. For tasks that were natively-sourced (such as TD) or were previously translated from English to Filipino (such as MT, QA), we filtered entries by text quality. Our evaluation of language quality was driven by acceptability judgments from native Filipino speakers [^4] Demographics: Working professionals, research faculty, and university students, aged between 21–30 years old. Native bilingual speakers and code-switchers of Filipino and English. Lived in the Philippines for a majority of their lives. in groups of three so that context and natural language variation are appropriately considered. Relying on standardized automated metrics would be less effective for languages with rich morphological and syntactic traits Rivera-Trigueros (2022) such as Filipino.
Annotation for SA task. We chose to repurpose the natively-sourced TD dataset due to the authenticity and diversity of registers in the dataset. Three native speakers annotated the Philippine election-related tweets Cabasag et al. (2019) dataset for the SA task with three sentiment polarities: positive, negative, and neutral. Inter-annotator agreement was calculated using Cohen’s kappa Cohen (1960), which is the difference between observed agreement and the probability of chance agreement over the probability that the raters did not agree by chance, and Krippendorff’s alpha Krippendorff (2018), which is 1 minus the ratio of observed disagreement by raters and expected chance disagreement). The resulting Cohen’s kappa of 0.8202 and Krippendorf’s alpha of 0.8268 indicate substantial agreement. For both TD and SA tasks, we chose to maintain the naturally-occurring language to capture the nuances of code-switched Taglish usage prevalent in social media in the Philippines.
Authoritative translation guidelines. For datasets that were adapted into Filipino, a critical aspect of dataset design was ensuring compliance with official guidelines set by the Komisyon sa Wikang Filipino (Commission on the Filipino Language) of the Philippines. In translation, we followed prescribed principles aimed at preserving the source material’s intent while adjusting for cultural and contextual relevance Almario (2016) and effective writing Almario (2014). Translations were performed by native Filipino speakers in the identified lingua franca, whenever appropriate, to maintain semantic equivalence without resorting to overly literal phrasing.
Native re-translation of previously translated datasets. We observed that the QA and MT datasets were of subpar quality and suffered from issues with inaccuracies and translationese Gellerstam (1986); Riley et al. (2020) such as the use of unnaturally formal language or inappropriate word choices. As such, these were re-translated to enhance contextual and linguistic accuracy.
Native translation of English datasets. For multilingual tasks that had no equivalent Filipino subsets (such as AS, CR, NLI, and PI), we generated initial translations using Helsinki NLP’s OPUS model. This model was selected over other off-the-shelf tools, such as Google Translate, as it demonstrated superior performance in terms of Filipino translation quality and accuracy as tested by the authors. These translations were then manually revised by native Filipino speakers to generate high-quality and fluent translations. Our native-speaker-driven evaluation process focused on cultural and linguistic relevance to ensure that the translated texts resonate with Filipino contexts.
3.4 Dataset Evaluation
Human evaluation for data quality. Each sample in each dataset underwent further quality assessment by teams of three raters against three binary criteria: completeness, fluency, and sensibility. Samples were assessed as complete only if they contained well-formed sentences, excluding standalone dependent clauses or fragmentary headers and titles lacking complete meaning. Fluency was evaluated on native-like constructions based on multiple factors including appropriate verb conjugations (e.g., avoiding constructions like “*Nagpapasalamatan ako sa iyo”, lit. “*I give thanks to you.”), natural word choices (avoiding awkward phrases like “*Si Pacquiao ang kamao ng bansa” instead of “Si Pacquiao ang pambansang kamao”, lit.: “Pacquiao is the nation’s fist”), and preference for KA sentence structure where contextually appropriate. Samples were assessed as sensible if relevant to the task, screening for confounding factors language capability (e.g., incorrect answers in QA mislabeled as correct or requiring re-translation).
Additional criteria for abstractive summarization. We employed additional criteria based on previous work Leong et al. (2023) to assess the appropriateness of the summary in reference to the content of the provided passage. We scored samples on binary criteria (completeness, fluency, and faithfulness) and a Likert-scaled (0-3) criteria (relevance, fluency, and coherence of the summary), with three independent raters per entry. For faithfulness, raters evaluated the factual consistency between the summary and source article, with particular attention to fact preservation and penalization of hallucinated content. For relevance, raters evaluated both coverage of essential information and information filtering, penalizing summaries containing redundancies or excess details. For fluency, raters assessed for proper formatting, appropriate capitalization, and grammatical construction, with higher scores awarded to summaries exhibiting natural Filipino language patterns. For coherence, raters assessed how effectively information flowed between sentences, with higher scores given to summaries that built logically from sentence to sentence to create a cohesive narrative about the topic.
3.5 Dataset Statistics
Competency Task Label # test samples
NLU PI True 200
False 200
QA span 100
SA Negative (Negatibo) 200
Neutral (Neutral) 200
Positive (Positibo) 200
TD Clean (Malinis) 200
Toxic (Mapoot) 200
NLR CR Cause (Sanhi) 200
Effect (Bunga) 200
NLI Contradiction 200
Entailment 200
Neutral 200
NLG AS summary 100
MT engtgl translation 600
tgleng translation 600
Total 3,800
Table 2: Class distribution per task in Batayan.
For all datasets except AS, only entries that were unanimously rated as demonstrating all 3/3 binary criteria were sampled for the final test splits. For AS, only entries that received a unanimous 3/3 binary criteria assessment and at least and average score of 2 for the scaled criteria were considered for inclusion, ensuring strict maintenance of factual integrity. In cases where initial sampling yielded insufficient qualifying entries, deficient samples were carefully corrected by authors through targeted improvements to grammar, translation of English passages, or other identified issues while maintaining authentic Filipino language patterns.
Joint agreement on our evaluation criteria is presented in Appendix C. We present joint agreement over traditional metrics such as Cohen’s kappa and Krippendorff’s alpha because metrics that use chance correction are less suitable for binary categorical tasks Powers (2012) and classification tasks in general Delgado and Tibau (2019); Feng and Zhao (2016).
Overall, Batayan provides 8 distinct tasks with 3,800 test instances. Table 2 provides the distribution of classes for each dataset, and more detailed statistics can be found in Appendix A. We also provided example splits composed of 5 entries that serve as exemplars for few-shot prompting.
4 Challenges with Developing a Filipino Benchmark
This section discusses issues in creating Batayan, shedding light on issues that future researchers may encounter in creating new Filipino datasets or adapting previous ones.
4.1 Issues with English-adapted and English-sourced Data
Creating more relevant summaries for XL-Sum. Previous work has noted that a significant portion of reference summaries in the XL-Sum dataset is highly abstractive, demonstrates factual errors, or contains information not mentioned in the provided articles Guo et al. (2022), putting to question its factuality and validity. To address this, we developed new Filipino summaries for each article. We ensured relevance and fluency by including only the most important information that can be directly lifted from the article using natural-sounding and grammatically-correct constructions.
Limitations in adapting vocabulary. Adapting the English tasks to Filipino revealed issues in maintaining the intelligibility of the text. Technical terms that exist in English were difficult to translate into Filipino, with several of them having no direct translations. For example, an instance from Belebele included the term “rule of thirds”, which describes a specific photography technique. Originally, the translation of this phrase in the Tagalog subset of Belebele was “tuntunin ng mga sangkatlo” (lit. “rule of thirds”), which does not make sense in Filipino and does not convey the intended meaning. For jargon such as this, we chose to keep the original English phrasing.
English idiomatic expressions were also prevalent and required a more flexible approach in adaptation to ensure the integrity of meaning. For example, one premise from Balanced COPA was, “The criminal turned himself in”, using the idiomatic expression “turned himself in” to mean “surrendered”. We chose to translate this into Filipino as, “Isinuko ng kriminal ang kanyang sarili” (lit.: “The criminal surrendered himself”).
Moreover, the presence of English words that have homonyms added to the complexity of translating English sentences to Filipino. For example, some sentences from PAWS contained the adjective “right”, which can refer to either the direction (in Filipino: “kanan”) or an assertion of the correctness of the object it is modifying (in Filipino: “tama”). In such cases, we inferred the most probable meaning of the words from the context of the entire passage and then identified the corresponding Filipino word to be incorporated in our Filipino translation.
(Dis)fluency in expert and machine translations. Our rigorous review of existing English datasets with machine and expert-guided translations in Filipino revealed weaknesses in their adaptation. The authors found that the automatic translations were unnatural, disfluent, and showed characteristics of translationese despite being grammatically correct. Notably, we observed that these initial translations demonstrated an unusual preference with using ay-inversion, and used Filipino terms that were synonymous to their English counterparts but pragmatically-awkward. Hence, we applied re-translation to ensure that the samples sounded more native and natural.
For example, one passage from the English subset of Belebele was, “The CCTV would certainly send a strong signal…” The original Tagalog (machine) translation was, “Ang CCTV ay tiyak na magpapadala ng malakas na hudyat” The term “magpapadala” here means “to send” in the sense of transporting a thing from one place to another, while the term “hudyat” implies a “starting sign”. It also uses the unnatural DKA construction or ay-inversion. Within the context of the original passage, a more fluent (human) translation would be, “Tiyak na maghahatid ang CCTV ng malakas na pahiwatig”, where “maghahatid” means “to bring about” and “pahiwatig” conveys a sense of “reminder” or “warning”, and the sentence follows the usual KA construction.
Inconsistencies in ay-inversions in translation construction. Styles in translation vary from person-to-person, depending on their valuations and reading of the original text Castagnoli (2020). Likewise, ay-inversion in the context of translation boils down to a stylistic choice of the translator.
While KA is argued to be the natural pattern of speech for Filipino speakers Yapan (2017), this valuation of natural and unnatural is debated. Magracia (2001) notes that while DKA is correct in “meaning, syntactic order, and grammaticality,” it is not the “natural” way of speaking for Filipinos, as it reflects the speaker’s language of thought—in this case, the influence of English. The same justification, in terms of source language for translated corpora, can be said for machine-translated text Visweswariah et al. (2011), though this remains an open question for Filipino.
On the other hand, other research argues that DKA is a necessary aspect for “discourse continuity” Bolata (2022). Hence, the more complex the sentence, the higher the chance of utilizing the ay-inversion Fox (1985). We observed this in some datasets such as NLI, which had longer, more context-rich sentences that required the use of DKA in contrast with the shorter, KA-ordered sentences found in the CR task.
These frameworks demonstrate the subjective nature of the selection of Filipino sentence structure. Hence, such inconsistencies may still be observed in Batayan despite being rigorously translated by native speakers.
4.2 Issues with Natively-Sourced Data
While the issues with most datasets in Batayan were related with fluency and naturalness, the Philippine election-related tweets Cabasag et al. (2019) dataset used for the SA and TD tasks presented unique challenges.
Incomplete entries. Due to the contextual nature of Filipino and social media communication, many samples lacked sufficient contextual information. To maintain dataset authenticity, we prioritized entries with adequate information for sentiment or toxicity analysis rather than complete removal. This was accomplished by selecting high-agreement samples where native speakers demonstrated consistent task performance.
Non-standard orthography. The dataset exhibited significant orthographic variation (spelling, capitalization, word boundaries, etc.) due to its natural language origins. We preserved this variation as it reflects authentic Filipino language use Ilao et al. (2012); Javier (2018); Caroro et al. (2020) relevant to toxicity detection and sentiment analysis tasks, aligning with our methodological principles. We maintained quality control by verifying that orthographical variations remained comprehensible to native readers.
Class imbalance. The dataset showed a disproportionate distribution of negative sentiment for certain political entities. We rebalanced the class distribution and prioritized high-agreement samples to mitigate individual annotator bias.
5 Evaluation
Batayan serves as the Filipino component of the SEA-HELM leaderboard. The evaluation framework of SEA-HELM follows prior systems such as HELM Liang et al. (2023).
5.1 Evaluation Design
Tasks. The selection of tasks in Batayan mentioned in Section 3 aims to comprehensively evaluate the Filipino language capabilities of LLMs. Each task is a collection of test and example instances composed of an input string, references, metadata, and a ground truth label.
Prompts. For Batayan, we developed prompt templates for each task written in Filipino. We ensured that the instructions for each task are consistent with the prompt template design already used in SEA-HELM. A comprehensive list of these prompt templates can be found in Appendix D.
During model evaluation, input prompts are constructed using evaluation instances and the corresponding prompt template. The default evaluation setting of Batayan for instruction-tuned models is zero-shot prompting, where model prompts do not include in-context input-label examples. Base pre-trained models without instruction-tuning are evaluated under a five-shot setting. Given the input prompts and decoding parameters (see Appendix LABEL:sec:experimental_setup), a model then generates output completions.
Metrics. We adopt multiple metrics to quantify the models’ performance on each task. For each metric, the model completion is treated as the prediction, while the instance label is used as the reference. For the NLU and NLR tasks, we report the macro F1 score. For machine translation (English→Filipino and Filipino→English), we use ChrF++ Popović (2017) and MetricX-24 using the metricx-24-hybrid-xxl-v2p6-bfloat16 model Juraska et al. (2024). For abstractive summarization, we report three metrics: BERTScore Zhang et al. (2020), ChrF++, and ROUGE-L F1 from the multilingual implementation of ROUGE Lin (2004) used in XL-Sum Hasan et al. (2021). Default package parameter settings are used, and performance scores are based on a single run of the benchmark.
5.2 Results
Figure 3: Model performance of selected models on NLU, NLR, and NLG tasks. NLU and NLR tasks are measured using macro F1 score, NLG tasks are measured using ChrF++, MetricX-24, BERTScore, and ROUGE-L F1. The unweighted average across tasks and scores is taken. Targeted Filipino instruction tuning allows Gemma-SEA-LION-v3-9B-IT to achieve competitive performance with significantly larger open-source and commercial models.
Our evaluation results include over 50 small ($\leq$11B), medium ($\leq$32B), and large ($\geq$32B) LLMs, including several commercial AI systems on the Batayan benchmark. The results underscore two principal findings: the substantial benefit of explicit Filipino language support and the scaling effects observed across model sizes. Figure 3 highlights several model performances, and the complete experiment results can be found in Appendix LABEL:sec:complete_experimental_results.
Filipino language support enhances task-specific performance. Tables LABEL:tab:small_nlu_nlrLABEL:tab:commercial_nlu_nlr in Appendix LABEL:sec:complete_experimental_results reveal that models with Filipino instruction tuning (such as the SEA-LION family of models) consistently outperform counterparts on culturally nuanced tasks. In particular, the Filipino-adapted model Gemma-SEA-LION-v3-9B-IT achieves an average macro F1 of 79.23 among small models, with a striking CR score of 92.75. In contrast, non-Filipino-specific models, such as many variants in the Meta Llama 3 series, exhibit very low or zero performance in CR, which emphasizes the challenge of capturing Filipino’s complex causality and morphology without explicit tuning.
On the NLG competency, tables LABEL:tab:small_nlgLABEL:tab:commercial_nlg in Appendix LABEL:sec:complete_experimental_results show that while ROUGE-L F1 scores show only modest differences between Filipino-tuned and general models, complementary evaluation using MetricX-24 provides a more sensitive measure of semantic fidelity. For instance, in engtgl translation, Filipino-supported systems like Gemma-SEA-LION-v3-9B-IT and Sailor2-8B-Chat achieve MetricX-24 scores exceeding 87, indicating superior preservation of Filipino-specific discourse and semantics. This observation shows that the apparent parity in ROUGE-L is only in part: when examined alongside semantic metrics, the benefit of dedicated language tuning is clear.
Scaling effects have clear trends beyond parameter similarity. Although many small models lie within a narrow 7B–9B parameter range, our broader evaluation—including medium and large models—illustrates a clear scaling effect. For example, among large models, Gemma-SEA-LION-v3-9B-IT attains an average macro F1 of 78.41 for NLU, macro F1 of 80.69 for NLR, and aggregated score of 60.35 in NLG, significantly outperforming similarly-tuned small models. This trend is evident even when models are compared within the same family: larger variants consistently demonstrate enhanced capability in handling both the nuanced understanding and generation challenges posed by Filipino text. The results thus suggest that increased model capacity—when combined with region-specific instruction tuning—enables a better representation of Filipino linguistic phenomena, including code-switching and agglutinative morphology.
Insights from commercial systems. In addition to open-source models, our evaluation also includes several commercial LLMs (Tables LABEL:tab:commercial_nlu_nlr and LABEL:tab:commercial_nlg). Commercial systems such as Gemini and GPT-4o variants exhibit competitive performance across both understanding and generation tasks. For instance, on NLU and NLR tasks, commercial models achieve average macro F1 scores ranging from 75.14 to 86.23. On the NLG competency, commercial systems also deliver robust results: Gemini and GPT-4o models maintain strong MetricX-24 scores above 88, indicative of high-quality semantic translation in both engtgl and tgleng directions.
While these commercial models benefit from large-scale pre-training and domain-specific optimization, our findings show that open-source, Filipino-tuned models achieve comparable—and in certain tasks, superior—performance. In particular, when cultural and linguistic nuances are critical, explicit regional instruction tuning offers a tangible advantage. The comparable performance of commercial systems reinforces that our dual strategy of targeted language support combined with model scaling is effective in academic settings as well as real-world applications.
Overall, our results provide strong evidence for a dual strategy in multilingual modeling: integrating region-specific instruction tuning with scalable architectures. Filipino-tuned models excel in tasks that require nuanced cultural and morphological sensitivity, and their performance improves markedly with increased capacity. Notably, open-source models fine-tuned with Filipino instructions are competitive with and even surpass state-of-the-art commercial systems on both understanding and reasoning tasks. These findings highlight that explicit Filipino language support not only enhances semantic fidelity and discourse preservation but also positions such models ideally for real-world, production-level applications in the Philippines.
6 Conclusion
In this paper, we introduced Batayan for holistically evaluating LLMs on a gamut of Filipino language tasks covering natural language understanding, reasoning, and generation. Our findings show that LLMs with explicit Filipino language support and fine-tuned on Filipino instructions demonstrate better performance on Batayan compared to other models.
As one of the major challenges in developing Batayan is maintaining the naturalness and fluency of the language, we plan to develop metrics and tools that can help discriminate from translationese and natural texts, inspired by prior research Lovenia et al. (2024); Riley et al. (2020).
Limitations
While our work provides a unique and comprehensive Filipino language benchmark, we also reported in previous sections the challenges and limitations in developing high-quality NLP resources for the under-represented language. Additionally, prosodic characteristics such as stress and sarcasm are not immediately obvious in the datasets utilized. We account for this by selecting only high-agreement samples. We also note that the domain of both SA and TD tasks (which were tweets surrounding political events), limits the distributions of these tasks, and as such we recommend utilizing other datasets that cover a wider set of domains and use cases.
While we recognize that datasets with larger amounts of samples are preferred to ensure generalizability, we maintain our position to prioritize consistent, high-quality, and representative samples. The thoroughness and rigor of our review, annotation, and translation processes were essential and maintaining the quality and reliability of our benchmark. We believe that our benchmark serves as a valuable starting point for future research and can certainly be expanded in subsequent iterations.
Further, given the breadth of models and tasks, a comprehensive evaluation incorporating error cases was not possible. We also recognize the importance of evaluation LLMs that specialize on specific tasks such as translation. We hope that a public release of the Batayan codebase and benchmark on a leaderboard can help researchers better characterize the weaknesses of models, conduct their own analysis on the Filipino language capabilities of LLMs, and experiment with specialized models on Batayan.
Ethical Considerations
The methodology used in this work was approved by an Institutional Review Board (NUS-IRB-2024-617). We have released the Batayan dataset as part of the SEA-HELM evaluation suite and leaderboard under the Creative Commons Attribution Share-Alike 4.0 (CC-BY-SA 4.0) license, respecting the licenses of the source datasets used in this study.
Due to the nature of the toxicity detection task, we note that the authors were exposed to offensive material. Nonetheless, they were encouraged to report inappropriate samples and were given the option to stop work if desired.
For the annotation of the sentiment analysis task, we involved a quality assurance team consisting of native Filipino speakers. The team, comprised of students from local universities in Singapore, were recruited through public advertisements. These stated the estimated work load and remuneration, which were consistent with university research guidelines and regulatory requirements.
We do not foresee negative social impacts from this paper. Our work introduces Filipino language resources that were reviewed by native Filipino speakers, paying due respect to local cultural sensitivities. We thus do not believe that our research will contribute to over-generalizations regarding Filipino culture.
Acknowledgments
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore.
The authors would like to thank National University of Singapore and AI Singapore, especially the AI Products and SEA-LION team, for their unwavering support with this endeavor.
The authors likewise thank the University of Cambridge, Department of Computer Science and Technology, and Magdalene College. David Africa’s work is supported by the Cambridge Trust and the Jardine Foundation.
Lastly, the authors would like to thank all the Filipino natives involved in this study for their time and valuable contributions.
Appendix A Overview of Tasks and Datasets
As discussed in Section 3 of this paper, the selection of the eight tasks are presented below. We also provide Table 3 describe the source datasets used for these tasks. The modifications and adaptations that were applied and the usage of the datasets for evaluation comply with their original intended uses.
Abstractive Summarization (AS). An LLM performing this task is given a paragraph and is expected to summarize the content in a sentence. The model is tested not only for identifying the salient points of the text, but also paraphrasing the content into a concise and coherent text. For this task, we use XL-Sum Hasan et al. (2021), a collection of annotated article-summary pairs.
Causal Reasoning (CR). This task requires the LLM to understand the relationship between events. In particular, the model is given a premise, a set of statements, and an instruction to determine which of the provided statements is the cause or effect of the premise. We employed Balanced COPA Kavumba et al. (2019), a dataset designed to evaluate commonsense causal reasoning with paired alternatives.
Machine Translation (MT). For this task, an LLM is given a text in one language, and is expected to provide the equivalent text translated into another language. In this study, we test for both English→Filipino and Filipino→English translation. This task leveraged the Filipino subset of FLORES 200 NLLB Team et al. (2022), which includes translations across numerous languages and domains.
Natural Language Inference (NLI). This is classification task where an LLM is provided two sentences (X and Y), and is expected to determine whether the sentences are related in one in the following ways: (a) X implies Y; (b) X contradicts Y; and (c) X neither implies nor contradicts Y. This task utilized XNLI Conneau et al. (2018), a dataset containing human-annotated examples for evaluating cross-lingual inference.
Paraphrase Identification (PI). This task requires an LLM to determine if two provided texts are paraphrased versions of each other; that is, whether both pieces of text convey the same idea. For this task, we used PAWS Zhang et al. (2019), which contains paraphrase and non-paraphrase pairs with high lexical overlap.
Question Answering (QA). Given a passage and a question, an LLM performing this task must provide a span from the passage that answers the question. In this study, QA utilized Belebele Bandarkar et al. (2024), a multiple-choice reading comprehension dataset designed to assess understanding of passages.
Toxicity Detection (TD) and Sentiment Analysis (SA). These two NLU tasks both require an LLM to analyze a natural language text. The TD task requires the model to identify whether hate speech and abusive language is used in the text, while the SA task requires the model to classify the text as either positive, negative, or neutral in sentiment polarity. Both tasks are derived from the Philippine election-related tweets dataset Cabasag et al. (2019), which provided a resource for toxicity in political discourse. These tweets were collected during the 2016 presidential campaign. This dataset was also relabeled for sentiment analysis by three native Filipino speakers. There is substantial agreement between the annotations (Cohen’s kappa of 0.8202, Krippendorf’s alpha of 0.8268).
Competency Task Dataset License
NLU PI PAWS Zhang et al. (2019) CC BY 4.0
QA Belebele Bandarkar et al. (2024) CC BY-NC 4.0
SA PH Elections Cabasag et al. (2019) Unknown
TD PH Elections Cabasag et al. (2019) Unknown
NLR CR Balanced COPA Kavumba et al. (2019) CC BY 4.0
NLI XNLI Conneau et al. (2018) CC BY-NC 4.0
NLG AS XL-Sum Hasan et al. (2021) CC BY-NC-SA 4.0
MT FLORES 200 NLLB Team et al. (2022) CC BY-SA 4.0
Table 3: License details for datasets used in Batayan.
Competency Task # test samples Avg. No. Words/Sample
NLU PI 400 22.50 (Sentence 1), 22.47 (Sentence 2)
QA 100 14.75 (Question), 4.26 (Choices)
SA 600 24.32
TD 400 20.73
NLR CR 400 7.24 (Premise), 6.05 (Choices)
NLI 600 23.30 (Sentence 1), 12.20 (Sentence 2)
NLG AS 100 120.8
MT (English Text) 600 21.42
MT (Tagalog Text) 600 25.05
Table 4: Summary of dataset statistics, including total rows and average words per text.
Table 4 presents quantitative statistics on dataset size and average word counts. The variation in text length reflects task-specific requirements: AS involves longer passages, while classification tasks (SA, TD, and PI) typically contain more concise text samples. The QA dataset, in particular, has relatively short questions and answer choices, mirroring real-world multiple-choice assessments. Further granularity is provided in Table 5, which breaks down sentence, word, and character-level statistics. This statistics shows structural differences between tasks; for instance, MT has sentence-level parallelism, while CR and NLI involve distinct premise-hypothesis or question-choice relationships.
Competency Task Component Avg. No. Sentences Avg. No. Words Avg. No. Characters
NLU PI Sentence 1 1.0 22.50 123.01
PI Sentence 2 1.0 22.47 122.99
QA Question 1.0 14.75 80.62
QA Choice 1 1.0 4.66 29.10
QA Choice 2 1.0 4.61 28.60
QA Choice 3 1.0 4.76 29.74
QA Choice 4 1.0 3.0 28.58
SA Text 2.21 24.32 120.44
TD Text 2.45 20.73 99.37
NLR CR Premise 1.0 7.24 37.55
CR Choice 1 1.0 6.05 30.91
CR Choice 2 1.0 6.05 30.62
NLI Sentence 1 1.0 23.30 127.88
NLI Sentence 2 1.0 12.20 67.04
NLG AS Text 5.43 120.8 666.18
MT English Text 1.12 21.42 113.07
MT Filipino Text 1.12 25.05 140.58
Table 5: Detailed dataset statistics, including sentence, word, and character counts.
Appendix B Dataset Examples
Competency Task Example
NLU PI Sentence 1: Wala pang isang taon matapos ang kaniyang nagdaang kasal, pinakasalan din ni Charlemagne si Desiderata at tinanggihan ang 13-anyos na Swabian na nagngangalang Hildegard.
Sentence 2: Pinakasalan ni Charlemagne si Desiderata wala pang isang taon pagkatapos ng kanyang kasal at isinawalang-bahala ang isang 13 taong gulang na Swabian na nagngangalang Hildegard.
Label: Paraprase (Paraphrase)
QA Text: Lumipat ang mga hukbong Coalition at Afghan sa lugar na iyon upang tiyaking ligtas ang lokasyon, at nagpadala ng iba pang eroplano ng koalisyon upang tumulong. Naganap ang pagbagsak sa mataas na bahagi ng kabundukan, at pinaniniwalaang resulta ng pagbabaril ng kalaban. Isang malaking hamon ang masamang panahon at malubak na daan sa paghahanap.
Question: Ano ang pinaniniwalaang dahilan ng pagbagsak?
Choices: (1) Pangit na daan (2) Masamang sunog (3) Mabundok na lupain (4) Masamang lagay ng panahon
Label: 1
SA Text: Pucha.. PURO DILAWAN DITO SA REDDIT AH HAHAHA.. AKALA NYO NAMAN ANG LAKI NA NG ACCOMPLISHMENT NYO DAHIL SA VIDEO NA YAN MGA GUNGGONG!! HAHAHA #DU30 PA RIN!! MGA TAE KAYO!
Label: Negatibo (Negative)
TD Text: Ayun, nag-Filipino rin si Poe, ang presidente ko! Kitang-kita ang kanyang platapormang maka-mahirap at maka-tao. #PiliPinasDebates2016
Label: Malinis (Clean)
NLR CR Text: Hindi tinanggap ang tsekeng ginawa ko.
Question: Sanhi (Cause)
Choices: (0) Walang laman ang bank account ko. (1) Tumaas ang aking suweldo.
Label: 0
NLI Sentence 1: Di mo ba natatandaan? Pupunta tayo ngayong araw sa birthday ni Tita Basia.
Sentence 2: Hindi kami pupunta sa birthday party ni Tita Basia ngayon.
Label: Contradiction
NLG AS Article: Kinumpirma noong nakaraang buwan na humawa na ang virus sa isang tupa sa isla matapos isilang nang patay at wala sa hugis ang limang kordero sa isang bukid. Sinabi ng state veterinary officer na malamang na ang birus ay sanhi ng windborne midges.
Summary: Kinumpirma ng mga pagsusuri noong nakaraang buwan na nahawahan ng Schmallenberg virus ang mga tupa sa isla.
MT Text: Former U.S. Speaker of the House Newt Gingrich came in second with 32 percent.
Translation: Pumangalawa ang dating Speaker of the House ng U.S. na si Newt Gingrich nang may 32 porsyento.
Table 6: Example instances for each task in Batayan.
Appendix C Inter-rater Agreement
Competency Task Dataset Criteria Joint agreement Rating
NLU PI PAWS Completeness 0.9550 -
Fluency 0.5875 -
Sensibility 0.8025 -
QA Belebele Completeness 0.9700 -
Fluency 0.8300 -
Sensibility 0.9500 -
SA PH Election Tweets Completeness 0.6217 -
Fluency 0.6767 -
Sensibility 0.7050 -
TD PH Election Tweets Completeness 0.8310 -
Fluency 0.9060 -
Sensibility 0.8680 -
NLR CR Balanced COPA Completeness 0.8975 -
Fluency 0.7275 -
Sensibility 0.9575 -
NLI XNLI Completeness 0.9716 -
Fluency 0.8433 -
Sensibility 0.9933 -
NLG AS XL-Sum Completeness 0.8800 -
Fluency 0.9600 -
Faithfulness 1.$0$
Relevance of summary - 1.99
Fluency of summary - 2.64
Coherence of summary - 2.56
MT FLORES 200 Completeness 0.7100 -
Fluency 0.9900 -
Sensibility 0.7750 -
Table 7: Joint agreement is calculated as the percentage of times all the raters unanimously agree that the samples fulfill the criteria. Rating is calculated as the average score assigned by the raters to the samples under the given Likert-scaled (0-3) criteria.
Appendix D Prompt Templates
Table 8: Performance of commercial LLMs on Batayan NLG tasks.
Task Filipino Prompt Template SEA-HELM English Prompt Template
PI Bibigyan ka ng dalawang pangungusap, SENTENCE_1 at SENTENCE_2. Tukuyin kung alin sa sumusunod na pahayag ang pinaka- angkop para sa SENTENCE_1 at SENTENCE_2. A: Paraprase ang SENTENCE_2 ng SENTENCE_1. B: Hindi paraprase ang SENTENCE_2 ng SENTENCE_1. Sumagot gamit ang sumusunod na format. Sagot: \$OPTION Palitan ang \$OPTION ng napiling sagot. Gumamit lang ng titik A o B sa sagot mo. {fewshot_examples} SENTENCE_1: ‘‘‘ {sentence1} ‘‘‘ SENTENCE_2: ‘‘‘ {sentence2} ‘‘‘ You will be given two sentences, SENTENCE_1 and SENTENCE_2. Determine which of the following statements applies to SENTENCE_1 and SENTENCE_2 the best. A: SENTENCE_2 is a paraphrase of SENTENCE_1. B: SENTENCE_2 is not a paraphrase of SENTENCE_1. Answer only using the following format: Answer: \$OPTION Replace \$OPTION with the selected option. Use the letters A or B only as the answer. {fewshot_examples} SENTENCE_1: ‘‘‘ {sentence1} ‘‘‘ SENTENCE_2: ‘‘‘ {sentence2} ‘‘‘
QA Bibigyan ka ng isang talata, isang tanong, at apat na pagpipiliang sagot. Sumagot base sa talata sa pamamagitan ng pagpili ng isa sa mga opsiyong ibinigay. Sumagot gamit ang sumusunod na format: Sagot: \$OPTION Palitan ang \$OPTION ng napiling sagot. Gumamit lang ng titik A, B, C, o D sa sagot mo.{fewshot_examples} Talata: ‘‘‘ {text} ‘‘‘ Tanong: {question} A: {choice1} B: {choice2} C: {choice3} D: {choice4} You will be given one paragraph, one question and four answer options. Based on the paragraph, answer by choosing one of the options given. Answer only using the following format: Answer: \$OPTION Replace \$OPTION with the selected option. Use the letter A, B, C, or D as the answer.{fewshot_examples} Paragraph: ‘‘‘ {text} ‘‘‘ Question: {question} A: {choice1} B: {choice2} C: {choice3} D: {choice4}
SA Ano ang sentimyento sa sumusunod na pangungusap? Sumagot gamit ng isa sa mga sumusunod na pagpipilian: Positibo, Negatibo, o Neutral. Sumagot gamit ang sumusunod na format: Sagot: \$OPTION Palitan ang \$OPTION ng napiling sagot. {fewshot_examples} Pangungusap: ‘‘‘ {text} ‘‘‘ What is the sentiment of the following sentence? Use one of the choices below: Positive, Negative, or Neutral. Answer only using the following format: Answer: \$OPTION Replace \$OPTION with the selected option. {fewshot_examples} Sentence: ‘‘‘ {text} ‘‘‘
TD Isa kang toxicity detector. Nangangahulugan ang mga sumusunod na label: Malinis: Walang toxicity. Mapoot: May toxicity at direktang pag-atake sa isang tao. Batay sa kahulugan ng mga label, iuri ang mga sumusunod na pangungusap gamit lang ng isang salita: Malinis o Mapoot. Sumagot gamit ang sumusunod na format: Sagot: \$LABEL Palitan ang \$LABEL ng napiling label. {fewshot_examples} Pangungusap: ‘‘‘ {text} ‘‘‘ You are a toxicity detector. The following labels are defined as follows: Clean: There is no toxicity. Toxic: There is toxicity and direct attacks towards a person. Based on the label definitions, classify the following sentence only using one word: Clean or Toxic. Answer only using the following format: Answer: \$LABEL Replace \$LABEL with the selected label. {fewshot_examples} Sentence: ‘‘‘ {text} ‘‘‘
CR Sumagot gamit ang sumusunod na format: Sagot: \$OPTION Palitan ang \$OPTION ng napiling sagot. Gumamit lang ng titik A or B sa sagot mo. {fewshot_examples} Batay sa ibibigay na sitwasyon, alin sa sumusunod na pagpipilian ang mas maaari na {sanhi/bunga}? Sitwasyon: ‘‘‘ {text} ‘‘‘ Piliin ang pinaka-angkop na sagot mula sa sumusunod na pagpipilian: A: {choice1} B: {choice2} Answer only using the following format: Answer: \$OPTION Replace \$OPTION with the selected option. Use only the letters A or B as the answer. {fewshot_examples} Based on the given situation, which of the following options is more likely to be the {cause/effect}? Situation: ‘‘‘ {text} ‘‘‘ Choose the best answer from the following options: A: {choice1} B: {choice2}
NLI Bibigyan ka ng dalawang pangungusap, SENTENCE_1 at SENTENCE_2. Tukuyin kung alin sa sumusunod na pahayag ang pinaka- angkop para sa SENTENCE_1 at SENTENCE_2. A: Kung totoo ang SENTENCE_1, dapat totoo din ang SENTENCE_2. B: Sumasalungat ang SENTENCE_1 sa SENTENCE_2. C: Kapag totoo ang SENTENCE_1, pwedeng totoo o hindi totoo ang SENTENCE_2. Sumagot gamit ang sumusunod na format. Sagot: \$OPTION Palitan ang \$OPTION ng napiling sagot. Gumamit lang ng titik A, B, o C sa sagot mo.{fewshot_examples} SENTENCE_1: ‘‘‘ {sentence1} ‘‘‘ SENTENCE_2: ‘‘‘ {sentence2} ‘‘‘ You will be given two sentences, SENTENCE_1 and SENTENCE_2. Determine which of the following statements applies to SENTENCE_1 and SENTENCE_2 the best. A: If SENTENCE_1 is true, SENTENCE_2 must be true. B: SENTENCE_1 contradicts SENTENCE_2. C: When SENTENCE_1 is true, SENTENCE_2 may or may not be true. Answer only using the following format: Answer: \$OPTION Replace \$OPTION with the selected option. Use the letters A, B or C only as the answer.{fewshot_examples} SENTENCE_1: ‘‘‘ {sentence1} ‘‘‘ SENTENCE_2: ‘‘‘ {sentence2} ‘‘‘
AS Ibuod ang sumusunod na artikulong Filipino sa isang talata na may isa o dalawang pangungusap. Sumagot gamit ang sumusunod na format: Buod: \$SUMMARY Palitan ang \$SUMMARY ng buod. {fewshot_examples} Artikulo: ‘‘‘ {text} ‘‘‘ Summarize the following {language} article into a paragraph with 1 or 2 sentences. Answer only using the following format: Summary: \$SUMMARY Replace \$SUMMARY with the summary. {fewshot_examples} Article: ‘‘‘ {text} ‘‘‘
MT Isalin ang sumusunod na teksto sa {language}. Sumagot gamit ang sumusunod na format: Salin: \$TRANSLATION Palitan ang \$TRANSLATION ng isinalin na teksto.{fewshot_examples} Teksto: ‘‘‘ {text} ‘‘‘ Translate the following text into {language}. Answer only using the following format: Translation: \$TRANSLATION Replace \$TRANSLATION with the translated text.{fewshot_examples} Text: ‘‘‘ {text} ‘‘‘
Appendix E Details of Models Evaluated
Models Source Link No. params
Base pre-trained models
aisingapore/Gemma-SEA-LION-v3-9B https://huggingface.co/aisingapore/Gemma-SEA-LION-v3-9B 9B
aisingapore/Llama-SEA-LION-v2-8B https://huggingface.co/aisingapore/Llama-SEA-LION-v2-8B 8B
aisingapore/Llama-SEA-LION-v3-8B https://huggingface.co/aisingapore/Llama-SEA-LION-v3-8B 8B
aisingapore/Llama-SEA-LION-v3-70B https://huggingface.co/aisingapore/Llama-SEA-LION-v3-70B 70B
google/gemma-2-9b https://huggingface.co/google/gemma-2-9b 9B
google/gemma-2-27b https://huggingface.co/google/gemma-2-27b 27B
google/gemma-3-12b-pt https://huggingface.co/google/gemma-3-12b-pt 12B
google/gemma-3-27b-pt https://huggingface.co/google/gemma-3-27b-pt 27B
meta-llama/Llama-3.1-8B https://huggingface.co/meta-llama/Llama-3.1-8B 8B
meta-llama/Llama-3.1-70B https://huggingface.co/meta-llama/Llama-3.1-70B 70B
Qwen/Qwen2-7B https://huggingface.co/Qwen/Qwen2-7B 7B
Qwen/Qwen2-72B https://huggingface.co/Qwen/Qwen2-72B 72B
Qwen/Qwen2.5-7B https://huggingface.co/Qwen/Qwen2.5-7B 7B
Qwen/Qwen2.5-14B https://huggingface.co/Qwen/Qwen2.5-14B 14B
Qwen/Qwen2.5-32B https://huggingface.co/Qwen/Qwen2.5-32B 32B
Qwen/Qwen2.5-72B https://huggingface.co/Qwen/Qwen2.5-72B 72B
sail/Sailor2-8B https://huggingface.co/sail/Sailor2-8B 8B
sail/Sailor2-20B https://huggingface.co/sail/Sailor2-20B 20B
SeaLLMs/SeaLLMs-v3-7B https://huggingface.co/SeaLLMs/SeaLLMs-v3-7B 7B
Instruction-tuned models
aisingapore/Gemma-SEA-LION-v3-9B-IT https://huggingface.co/aisingapore/Gemma-SEA-LION-v3-9B-IT 9B
aisingapore/Llama-SEA-LION-v2-8B-IT https://huggingface.co/aisingapore/Llama-SEA-LION-v2-8B-IT 8B
aisingapore/Llama-SEA-LION-v3-8B-IT https://huggingface.co/aisingapore/Llama-SEA-LION-v3-8B-IT 8B
aisingapore/Llama-SEA-LION-v3-70B-IT https://huggingface.co/aisingapore/Llama-SEA-LION-v3-70B-IT 70B
CohereForAI/aya-23-8B https://huggingface.co/CohereForAI/aya-23-8B 8B
CohereForAI/aya-23-35B https://huggingface.co/CohereForAI/aya-23-35B 35B
CohereForAI/aya-expanse-8b https://huggingface.co/CohereForAI/aya-expanse-8b 8B
CohereForAI/aya-expanse-32b https://huggingface.co/CohereForAI/aya-expanse-32b 32B
deepseek-ai/DeepSeek-V3 https://huggingface.co/deepseek-ai/DeepSeek-V3 671B
google/gemma-2-9b-it https://huggingface.co/google/gemma-2-9b-it 9B
google/gemma-2-27b-it https://huggingface.co/google/gemma-2-27b-it 27B
google/gemma-3-12b-it https://huggingface.co/google/gemma-3-12b-it 12B
google/gemma-3-27b-it https://huggingface.co/google/gemma-3-27b-it 27B
MERaLiON/Llama-3-MERaLiON-8B-Instruct https://huggingface.co/MERaLiON/Llama-3-MERaLiON-8B-Instruct 8B
meta-llama/Llama-3.1-8B-Instruct https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct 8B
meta-llama/Llama-3.1-70B-Instruct https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct 70B
meta-llama/Llama-3.3-70B-Instruct https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct 70B
Qwen/Qwen2-7B-Instruct https://huggingface.co/Qwen/Qwen2-7B-Instruct 7B
Qwen/Qwen2-72B-Instruct https://huggingface.co/Qwen/Qwen2-72B-Instruct 72B
Qwen/Qwen2.5-7B-Instruct https://huggingface.co/Qwen/Qwen2.5-7B-Instruct 7B
Qwen/Qwen2.5-14B-Instruct https://huggingface.co/Qwen/Qwen2.5-14B-Instruct 14B
Qwen/Qwen2.5-32B-Instruct https://huggingface.co/Qwen/Qwen2.5-32B-Instruct 32B
Qwen/Qwen2.5-72B-Instruct https://huggingface.co/Qwen/Qwen2.5-72B-Instruct 72B
sail/Sailor2-8B-Chat https://huggingface.co/sail/Sailor2-8B-Chat 8B
sail/Sailor2-20B-Chat https://huggingface.co/sail/Sailor2-20B-Chat 20B
SeaLLMs/SeaLLMs-v3-7B-Chat https://huggingface.co/SeaLLMs/SeaLLMs-v3-7B-Chat 7B
Tower-Babel/Babel-9B-Chat https://huggingface.co/Tower-Babel/Babel-9B-Chat 9B
Tower-Babel/Babel-83B-Chat https://huggingface.co/Tower-Babel/Babel-83B-Chat 83B
Reasoning models
deepseek-ai/DeepSeek-R1 https://huggingface.co/deepseek-ai/DeepSeek-R1 671B
Commercial models
google/gemini-1.5-flash-002 https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions -
google/gemini-1.5-pro-002 https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions -
google/gemini-2.0-flash-001 https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions -
google/gemini-2.0-flash-lite-001 https://cloud.google.com/vertex-ai/generative-ai/docs/learn/model-versions -
openai/gpt-4o-mini-2024-07-18 https://platform.openai.com/docs/models/gpt-4o-mini -
openai/gpt-4o-2024-08-06 https://platform.openai.com/docs/models/gpt-4o -
Model Reference Filipino IT Rationale for selection
Aya 23 and Expanse Aryabumi et al. (2024); Dang et al. (2024) LLMs with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Babel Zhao et al. (2025) LLMs with multilingual support including Filipino. This model was pre-trained and fine-tuned on Filipino tokens.
DeepSeek V3 and R1 DeepSeek-AI (2025) LLMs with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Gemma 2 and 3 Team et al. (2024b, 2025) LLMs with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Gemini 1.5 and 2 Team et al. (2024a) Commercial AI systems with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
GPT-4o and 4o-mini OpenAI et al. (2024) Commercial AI systems with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Llama 3 Grattafiori et al. (2024) LLMs with multilingual support. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
MERaLiON Huang et al. (2025) LLM with support for English, Chinese, and Indonesian. Llama 3 model was continue pre-trained on 120B English, Chinese, and Indonesian tokens. It is unknown if these models were explicitly fine-tuned on Filipino instructions.
Qwen 2 and 2.5 Yang et al. (2024); Qwen et al. (2025) LLMs with multilingual support, including Tagalog.
Sailor 2 Dou et al. (2025) LLMs with support for Southeast Asian languages. Qwen 2.5 7B model was continue pre-trained on 400B SEA tokens, and was specifically fine-tuned on Filipino instructions.
SEA-LION v2 and v3 Ng et al. (2025) LLMs with support for Southeast Asian languages. Llama 3 and Gemma 2 series models were continue pre-trained on 200B tokens across 11 Southeast Asian languages, and was specifically fine-tuned on Filipino instructions.
SeaLLMs v3 Zhang et al. (2024) LLMs with support for Southeast Asian languages. These models were pre-trained and fine-tuned on Filipino tokens.
Appendix F Experimental Setup
The default evaluation setting for Batayan is zero-shot prompting, however base pre-trained models with no instruction-tuning are evaluated on a five-shot setting. Performance scores are calculated based on a single run. Table LABEL:tab:inference_parameters details the decoding parameters used in the experiments.
Competency Task max_tokens temperature top_p top_k repetition_penalty
NLU PI 32 0.0 1.0 1.0 1.0
QA 32 0.0 1.0 1.0 1.0
SA 32 0.0 1.0 1.0 1.0
TD 32 0.0 1.0 1.0 1.0
NLR CR 32 0.0 1.0 1.0 1.0
NLI 32 0.0 1.0 1.0 1.0
NLG AS 512 0.3 1.0 1.0 1.0
MT 256 0.0 1.0 1.0 1.0
Appendix G Complete Experimental Results
Model PI QA SA TD NLU Avg. CR NLI NLR Avg.
Base pre-trained models (five-shot)
aisingapore/Gemma-SEA-LION-v3-9B 82.40 89.02 67.60 51.54 72.64 0.00a 63.93 31.97
aisingapore/Llama-SEA-LION-v2-8B 71.21 69.54 48.86 52.89 60.62 0.00a 28.99 14.49
aisingapore/Llama-SEA-LION-v3-8B 71.96 71.07 38.58 66.44 62.01 0.00a 48.59 24.30
google/gemma-2-9b 70.64 85.79 51.09 74.26 70.44 0.00a 46.04 23.02
meta-llama/Llama-3.1-8B 60.31 71.91 50.16 58.51 60.22 0.00a 26.82 13.41
Qwen/Qwen2-7B 84.69 74.12 72.11 65.48 74.10 0.00a 45.85 22.92
Qwen/Qwen2.5-7B 84.48 69.89 72.22 64.71 72.82 0.00a 41.11 20.55
sail/Sailor2-8B 79.44 75.93 70.25 67.64 73.32 0.00a 43.84 21.92
SeaLLMs/SeaLLMs-v3-7B 84.95 74.05 71.12 63.54 73.41 0.00a 41.46 20.73
Instruction-tuned models (zero-shot)
aisingapore/Gemma-SEA-LION-v3-9B-IT 82.00 82.80 75.99 72.85 78.41 92.75 68.64 80.69
aisingapore/Llama-SEA-LION-v2-8B-IT 68.50 45.41 49.81 58.04 55.44 50.23 37.63 43.93
aisingapore/Llama-SEA-LION-v3-8B-IT 74.97 80.69 51.24 60.57 66.87 79.44 65.87 72.65
CohereForAI/aya-23-8B 24.68 33.50 33.49 32.70 31.09 28.12 16.67 22.39
CohereForAI/aya-expanse-8b 32.65 2.97a 57.44 34.39 31.86 1.96a 25.27 13.61
google/gemma-2-9b-it 82.74 83.96 70.80 64.48 75.50 91.25 60.24 75.74
MERaLiON/Llama-3-MERaLiON-8B-Instruct 23.47 64.90 35.16 27.64 37.79 28.97 23.35 26.16
meta-llama/Llama-3.1-8B-Instruct 61.38 74.90 52.55 50.00 59.71 72.66 56.28 64.47
Qwen/Qwen2-7B-Instruct 73.08 56.01 52.63 26.10 51.96 43.99 32.35 38.17
Qwen/Qwen2.5-7B-Instruct 85.22 69.07 67.82 60.84 70.74 46.15 57.80 51.98
sail/Sailor2-8B-Chat 73.60 59.48 48.67 47.24 57.25 57.52 61.21 59.36
SeaLLMs/SeaLLMs-v3-7B-Chat 37.85 49.72 70.63 61.48 54.92 71.82 20.72 46.27
Tower-Babel/Babel-9B-Chat 78.76 61.68 72.52 56.74 67.42 83.00 26.89 54.95
  • a
    Could not follow answer instructions.
Model PI QA SA TD NLU Avg. CR NLI NLR Avg.
Base pre-trained models (five-shot)
google/gemma-2-27b 72.63 86.04 69.60 75.95 76.05 0.00a 56.21 28.11
google/gemma-3-12b-pt 71.42 86.02 79.30 51.60 72.08 0.00a 59.47 29.74
google/gemma-3-27b-pt 59.29 84.83 69.84 52.34 66.58 0.00a 55.13 27.56
Qwen/Qwen2.5-14B 89.25 82.03 68.36 62.89 75.63 0.00a 53.10 26.55
Qwen/Qwen2.5-32B 87.69 84.98 67.71 66.62 76.75 0.00a 55.09 27.54
sail/Sailor2-20B 85.46 84.94 68.47 48.83 71.92 0.00a 63.09 31.55
Instruction-tuned models (zero-shot)
CohereForAI/aya-expanse-32b 76.39 54.82 63.25 74.70 67.29 55.46 66.57 61.01
google/gemma-3-12b-it 82.27 83.94 62.10 80.50 77.20 93.75 70.64 82.19
google/gemma-2-27b-it 70.68 69.18 70.39 80.36 72.65 67.76 80.25 73.04
google/gemma-3-27b-it 70.20 87.91 67.77 81.74 76.91 94.75 67.73 81.24
Qwen/Qwen2.5-14B-Instruct 63.75 64.64 72.34 72.73 68.37 53.47 43.83 48.65
Qwen/Qwen2.5-32B-Instruct 83.70 84.97 74.60 75.50 79.69 85.98 60.58 73.28
sail/Sailor2-20B-Chat 39.17 13.81 54.41 74.03 45.36 22.98 47.88 35.43
  • a
    Could not follow answer instructions.
Model PI QA SA TD NLU Avg. CR NLI NLR Avg.
Base pre-trained models (five-shot)
aisingapore/Llama-SEA-LION-v3-70B 83.48 87.02 45.89 77.49 73.47 0.00a 79.62 39.81
meta-llama/Llama-3.1-70B 83.49 86.05 59.72 51.70 70.24 0.00a 55.78 27.89
Qwen/Qwen2-72B 90.25 88.86 67.88 77.51 81.13 0.00a 69.63 34.81
Qwen/Qwen2.5-72B 86.89 86.79 74.73 79.61 82.01 0.00a 67.58 33.79
Instruction-tuned models (zero-shot)
aisingapore/Llama-SEA-LION-v3-70B-IT 84.07 87.01 68.79 76.99 79.21 94.75 75.34 85.04
CohereForAI/aya-23-35B 81.83 61.46 61.68 53.06 64.51 74.16 40.54 57.35
deepseek-ai/DeepSeek-V3 83.07 90.00 81.78 76.21 82.77 94.25 63.25 78.75
meta-llama/Llama-3.1-70B-Instruct 83.57 88.13 67.29 63.80 75.70 93.49 72.62 83.06
meta-llama/Llama-3.3-70B-Instruct 82.62 86.96 66.12 77.40 78.28 94.25 72.97 83.61
Qwen/Qwen2-72B-Instruct 85.37 86.10 76.16 76.38 81.00 58.54 68.39 63.47
Qwen/Qwen2.5-72B-Instruct 80.40 84.00 75.32 73.44 78.29 91.75 63.95 77.85
Tower-Babel/Babel-83B-Chat 85.06 33.12 73.20 81.24 68.15 49.62 68.14 58.88
Reasoning models (zero-shot)
deepseek-ai/DeepSeek-R1 79.84 87.93 71.47 71.47 77.68 97.00 71.06 84.03
  • a
    Could not follow answer instructions.
Model PI QA SA TD NLU Avg. CR NLI NLR Avg.
google/gemini-1.5-flash-002 85.71 84.98 63.81 66.04 75.14 94.24 69.19 81.72
google/gemini-1.5-pro-002 85.62 89.95 69.48 79.27 81.08 95.74 76.71 86.23
google/gemini-2.0-flash-001 83.98 88.05 73.93 80.49 81.61 97.74 70.47 84.11
google/gemini-2.0-flash-lite-001 85.02 85.97 64.33 78.94 78.57 96.49 70.20 83.35
openai/gpt-4o-mini-2024-07-18 86.11 85.04 70.77 70.77 78.17 92.99 68.18 80.59
openai/gpt-4o-2024-08-06 88.21 71.88 74.78 74.78 77.41 97.25 73.33 85.29
Models AS MT (engtgl) MT (tgleng) NLG
BERTScore ChrF++ ROUGE-L ChrF++ MetricX-24 ChrF++ MetricX-24 Avg.
Base pre-trained models (five-shot)
aisingapore/Gemma-SEA-LION-v3-9B 77.01 34.19 33.71 58.90 82.75 56.63 91.15 62.05
aisingapore/Llama-SEA-LION-v2-8B 75.86 27.17 30.88 48.45 62.74 58.17 87.10 55.77
aisingapore/Llama-SEA-LION-v3-8B 75.73 32.41 30.56 53.10 72.41 57.70 88.91 58.69
google/gemma-2-9b 76.60 31.72 31.69 59.06 82.01 63.27 91.16 62.22
meta-llama/Llama-3.1-8B 75.80 29.93 30.13 51.94 71.10 58.44 88.77 58.01
Qwen/Qwen2-7B 73.84 27.64 25.71 42.61 47.37 53.60 82.13 50.41
Qwen/Qwen2.5-7B 73.26 26.71 25.04 38.89 40.28 42.59 81.58 46.91
sail/Sailor2-8B 77.05 30.69 34.50 62.01 86.30 63.41 91.05 63.57
SeaLLMs/SeaLLMs-v3-7B 73.78 28.37 26.57 43.45 47.79 44.54 83.74 49.75
Instruction-tuned models (zero-shot)
aisingapore/Gemma-SEA-LION-v3-9B-IT 73.00 34.16 20.99 57.79 87.24 58.63 90.56 60.35
aisingapore/Llama-SEA-LION-v2-8B-IT 73.39 32.91 24.14 50.90 64.62 48.49 55.04 49.93
aisingapore/Llama-SEA-LION-v3-8B-IT 73.34 33.80 23.43 57.14 82.67 58.33 87.69 59.49
CohereForAI/aya-23-8B 67.37 24.01 11.76 15.61 9.68 46.20 58.59 33.32
CohereForAI/aya-expanse-8b 59.48 6.69 1.64 39.30 37.47 49.81 78.88 39.04
google/gemma-2-9b-it 72.73 34.19 21.81 56.80 84.23 50.97 76.62 56.76
MERaLiON/Llama-3-MERaLiON-8B-Instruct 74.08 31.79 25.68 49.77 64.85 41.66 60.14 49.71
meta-llama/Llama-3.1-8B-Instruct 74.24 34.27 26.94 44.79 51.76 46.33 70.65 49.85
Qwen/Qwen2-7B-Instruct 69.51 26.09 13.58 42.22 42.86 45.13 50.01 41.34
Qwen/Qwen2.5-7B-Instruct 70.61 28.89 15.86 38.99 34.57 49.02 76.36 44.91
sail/Sailor2-8B-Chat 71.82 31.83 19.75 54.04 80.93 35.41 19.66 44.78
SeaLLMs/SeaLLMs-v3-7B-Chat 70.34 27.84 16.49 46.36 61.07 43.27 60.37 46.53
Tower-Babel/Babel-9B-Chat 71.61 31.55 18.93 49.97 72.50 48.69 75.69 52.70
Models AS MT (engtgl) MT (tgleng) NLG
BERTScore ChrF++ ROUGE-L ChrF++ MetricX-24 ChrF++ MetricX-24 Avg.
Base pre-trained models (five-shot)
google/gemma-2-27b 77.27 33.71 33.50 61.16 85.56 64.77 91.82 63.97
google/gemma-3-12b-pt 77.42 32.12 34.32 61.66 85.52 64.69 92.08 63.97
google/gemma-3-27b-pt 77.70 35.53 35.36 62.89 87.06 64.98 92.27 65.11
Qwen/Qwen2.5-14B 74.39 30.34 27.87 49.01 62.69 59.41 88.63 56.05
Qwen/Qwen2.5-32B 75.13 29.73 29.31 50.94 65.06 60.18 89.34 57.10
sail/Sailor2-20B 76.73 32.41 33.47 61.75 86.30 63.88 91.73 63.75
Instruction-tuned models (zero-shot)
CohereForAI/aya-expanse-32b 71.18 30.12 16.85 49.14 62.76 57.35 85.05 53.21
google/gemma-2-27b-it 73.41 35.11 23.21 59.51 88.19 61.87 91.68 61.86
google/gemma-3-12b-it 72.10 32.80 19.43 60.09 89.39 61.85 91.61 61.04
google/gemma-3-27b-it 72.25 33.26 19.55 61.02 90.71 62.09 92.18 61.58
Qwen/Qwen2.5-14B-Instruct 71.92 31.85 19.24 44.78 54.51 58.01 86.08 52.34
Qwen/Qwen2.5-32B-Instruct 72.47 33.43 21.22 47.47 59.47 58.77 86.90 54.25
sail/Sailor2-20B-Chat 71.47 31.20 18.41 49.45 72.51 49.52 72.79 52.19
Models AS MT (engtgl) MT (tgleng) NLG
BERTScore ChrF++ ROUGE-L ChrF++ MetricX-24 ChrF++ MetricX-24 Avg.
Base pre-trained models (five-shot)
aisingapore/Llama-SEA-LION-v3-70B 76.87 35.49 33.66 60.70 85.28 63.39 91.52 63.85
meta-llama/Llama-3.1-70B 76.80 33.98 33.54 60.20 84.32 65.80 91.85 63.79
Qwen/Qwen2-72B 76.81 34.15 33.24 54.75 74.84 65.13 90.87 61.40
Qwen/Qwen2.5-72B 76.41 34.12 31.96 54.69 74.92 64.50 90.85 61.06
Instruction-tuned models (zero-shot)
aisingapore/Llama-SEA-LION-v3-70B-IT 73.25 34.82 23.81 61.02 86.24 63.11 91.28 61.93
CohereForAI/aya-23-35B 72.81 29.94 22.08 46.00 56.79 49.54 72.45 49.94
deepseek-ai/DeepSeek-V3 73.19 35.16 22.99 61.61 89.21 63.06 92.28 62.50
meta-llama/Llama-3.1-70B-Instruct 74.76 35.56 27.94 60.32 84.73 62.78 91.20 62.47
meta-llama/Llama-3.3-70B-Instruct 72.90 34.03 22.63 58.98 83.05 61.24 90.72 60.51
Qwen/Qwen2-72B-Instruct 67.05 25.08 11.38 52.84 70.63 60.05 88.73 53.68
Qwen/Qwen2.5-72B-Instruct 71.62 31.52 18.02 52.01 72.30 64.33 90.02 57.12
Tower-Babel/Babel-83B-Chat 67.72 20.58 13.35 53.29 76.61 51.56 85.42 52.65
Reasoning models (zero-shot)
deepseek-ai/DeepSeek-R1 71.97 32.70 19.59 59.15 90.91 61.01 92.30 61.09
Models AS MT (engtgl) MT (tgleng) NLG
BERTScore ChrF++ ROUGE-L ChrF++ MetricX-24 ChrF++ MetricX-24 Avg.
google/gemini-1.5-flash-002 72.44 34.00 21.20 61.01 90.11 65.23 92.04 62.29
google/gemini-1.5-pro-002 73.62 36.16 24.00 61.07 90.90 64.14 91.30 63.03
google/gemini-2.0-flash-001 72.99 34.85 21.96 63.88 90.87 65.64 92.53 63.25
google/gemini-2.0-flash-lite-001 72.83 34.80 21.87 63.45 90.06 65.32 91.51 62.83
openai/gpt-4o-mini-2024-07-18 72.70 33.97 21.67 62.46 88.59 63.12 91.68 62.03
openai/gpt-4o-2024-08-06 72.83 34.36 21.55 63.86 90.19 64.89 92.51 62.88